OpenStudy (usukidoll):

I need my memory refreshed on this... Which of these sets are linearly dependent on the whole line? a. x-3, 6-2x, e^x b. x+1 2x-3, 3x+4 cos^2x cos2x, sin^2x d. 1, e^x, xe^x

OpenStudy (usukidoll):

I can see that a and b are all P1 vector spaces because of the degree.

OpenStudy (usukidoll):

and I remember that I have to put it like this c1(x-3)+c2(6-2x)+c3(e^x) = 0x + 0

OpenStudy (usukidoll):

so xc1-3c1+6c2-2xc2+c3e^x = 0 and I want to put it in an augmented matrix format

OpenStudy (usukidoll):

so that would be if I rearrange this xc1 -2xc2 0 -3c1+6c2 0 e^x c3 = 0